# A Comprehensive Framework to Operationalize Social Stereotypes for Responsible AI Evaluations

### v2

@@ 7 Limitations
While our framework captures various aspects of stereotyping by drawing from social psychology and NLP, we acknowledge its potential limitations. First, our goal is for the framework improve stereotype evaluation and mitigation in LLMs. This inherent subjectivity in interpreting the application can limit the generalizability of the framework to other NLP tasks. Second, while out framework emphasizes the essential role of “context” in shaping stereotypes, we recognize that context is inherently multifaceted and dynamic, encompassing a vast array of factors, including but not limited to social norms, historic events, individual experiences, and power dynamics. Due to this complexity, any attempt to model the context is inevitably meant to fail. Instead, we encourage researchers to explicitly consider and document the relevant contextual factors in their efforts, even if the those factors expand beyond the specific elements included in current framework. Ongoing critical engagement and reflection is highly necessary for social and historical grounding of stereotype evaluations.

### v3

@@ 7 Limitations
While our framework captures various aspects of stereotyping by drawing from social psychology and NLP, we acknowledge its potential limitations. First, our goal is for the framework to improve stereotype evaluation and mitigation in LLMs. This inherent focus on model-centric applications and the subjectivity in interpreting the application can limit the generalizability of the framework to other NLP tasks. Second, while our framework emphasizes the essential role of Context in shaping stereotypes, we recognize that context is inherently multifaceted and dynamic, encompassing a vast array of factors, including but not limited to social norms, historic events, individual experiences, and power dynamics. Due to this complexity, any attempt to model the context is inevitably incomplete. Instead, we encourage researchers to explicitly consider and document the relevant contextual factors in their efforts, even if those factors expand beyond the specific elements included in the current framework. Moreover, several studies in NLP tend to the linguistic context in which stereotypes are expressed and explore nuanced communication elements such as linguistic modalities, reasons, motivations, sarcasm, and parody as they co-occur with stereotyping language. A focused linguistic effort is essential for incorporating such linguistic factors with the core aspects of stereotypes discussed in this paper. Therefore, ongoing critical engagement and reflection is highly necessary for linguistic, social and historical grounding of stereotype evaluations.
Appendix A Glossary
Stereotype
— A cognitive generalization about a specific social group, often consisting of widely shared beliefs and assumed traits associated with its members.
Categorizing
— The fundamental cognitive process of grouping objects, events, or people into categories, which is essential to the formation of stereotypes.
Intersectionality
— The concept that individuals belong to multiple social groups simultaneously, and that stereotypes targeting these intersectional identities create unique forms of bias beyond those of the component groups.
Stereotype Content Model (SCM)
— A foundational social psychological framework that posits group stereotypes vary along two primary, universal dimensions: Warmth and Competence.
Agency-Beliefs-Communion Model (ABC)
— A theoretical extension of the SCM that adds Beliefs as a third dimension, alongside refined aspects of Agency (Competence) and Communion (Warmth).
Warmth
— The SCM dimension that captures perceived good or ill intent, reflecting traits like friendliness, sincerity, and morality.
Competence
— The SCM dimension that captures perceived capability or status, reflecting traits like intelligence, skill, and agency.
Perceiver
— The individual, group, or section of society that holds and applies a specific stereotypical belief about the target group.